Solve
10x^4-9x^2+2=0
a the solution is x=
b there is no solution

Answer by sudhanshu_kmr(1152)


10x^4-9x^2+2=0
=> 10x^4-5x^2-4x^2 + 2=0
=> 5x^2(2x^2-1)-2(2x^2-1)= 0
=> (5x^2 -2)(2x^2 -1) =0

x = root(2/5)
= - root(2/5)
= root(1/2)
= - root(1/2)